OpenStudy (anonymous):

x+2 / -x-6 >0 write answer using interval notation.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{x+2 }{-x-6}>0 }\) right?

OpenStudy (solomonzelman):

for the fraction to be positive both top and bottom have to be negative or both positive.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{x+2 }{-x-6}>0 }\) \(\large\color{black}{ \displaystyle \frac{x+2 }{-(x+6)}>0 }\) from anywhere less than -6, the bottom is positive and the top is negative anywhere more than -2 the bottom is negative and the top is positive.

OpenStudy (solomonzelman):

So where are both top and bottom negative or both positive?

OpenStudy (anonymous):

what? i not exactly sure what ur saying

OpenStudy (anonymous):

*Im